OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again? 0.6 sec 0.8 sec 1.6 secs 2.8 secs

OpenStudy (anonymous):

Please help if you know! :D

OpenStudy (anonymous):

Oh, ok, thx. :)

OpenStudy (anonymous):

we have a quadratic equation formula, we have to use. :)

OpenStudy (anonymous):

Do you have a quadratic equation for the height of the ball at some time "t"? Like a function h(t) = ...

OpenStudy (anonymous):

Hmmm...let me see. :)

OpenStudy (anonymous):

I don't know, but I am preety sure you are right.

OpenStudy (anonymous):

What class is this?

OpenStudy (anonymous):

Algebra 1. :) I just looked up on Google, and they said the answer was B. But I am not sure. :)

OpenStudy (anonymous):

I would have said (c)... here's why: acceleration of gravity = 32.174 feet per second squared velocity 1 = 25 feet per second velocity 2 = -25 feet per second (speed of ball when it hits the ground is the same as when it was kicked, but opposite direction) so time = (25 - (-25)) / 32.174 = 50 / 32.174 = 1.554 seconds

OpenStudy (anonymous):

I'm a little surprised they gave you this question in algebra 1. If they had given you the formula for the balls height as something like: h(t) = At^2 + Bt + C Then you could ask and answer, if the height is zero when it hits the ground, then what is the time, t? 0 = At^2 + Bt + C Then you could use the quadratic formula.

OpenStudy (anonymous):

But requires a height formula... you could derive that formula from the initial velocity, but that might require some calculus :)

OpenStudy (anonymous):

Oh, wow, I do not know what to put!

OpenStudy (anonymous):

Thx for your help!! :)

OpenStudy (anonymous):

I might have messed it up... I was working from memory, but 0.8 could be right.

OpenStudy (anonymous):

Yeah, I think I messed it up above. The equation is h(t) = at^2 + vt + d where a = acceleration of gravity = - 32.174 ft/sec^2 v = 25 feet/sec d = starting height = 0 and h(t) = 0 when the ball hits the ground. So, 0 = - 32.174t^2 + 25t + 0 You can use the quadratic formula on that if you want, or you can solve like this: 0 = - 32.174t^2 + 25t 0 = t ( -32.174t + 25) So, one solution of that is t = 0, corresponding to the initial time when the ball is kicked. The other time is: 25 = 32.174t t = 25/32.174 = 0.777 seconds or approximately 0.8 seconds.

OpenStudy (anonymous):

OMG, thx! That is what Google said. :)

OpenStudy (anonymous):

Yes, but I've found other explanations that said what I said first. I'm confusing myself...

OpenStudy (anonymous):

LOL, Algebra is confusing. :)

OpenStudy (anonymous):

I would go with 0.8, but this is better taught as a physics question. Sorry I wasn't more help!

OpenStudy (anonymous):

it s okay! I will go with that, and let youknow if it was right. :)